OpenStudy (anonymous):

Can someone check this? Solve the compound inequality. 4x+15>=-9 and 8x-6<=34 a.x>=6 and x<=4 b.x>=-6 and x<=-5 c.x>=-6 and x<=5 d.x>=6 and x<=5

OpenStudy (anonymous):

My choice is C.

rishavraj (rishavraj):

yup :))

OpenStudy (anonymous):

Would you mind checking another one? @rishavraj

rishavraj (rishavraj):

okkk

OpenStudy (anonymous):

This one? @rishavraj

rishavraj (rishavraj):

why +10/13 ???

rishavraj (rishavraj):

i mean why positive one??

rishavraj (rishavraj):

see 13y = 12x - 10 proceed

OpenStudy (anonymous):

I was thinking that the slope was correct and the y intercept as well

rishavraj (rishavraj):

see y = mx + c m - slope c - y intercept

rishavraj (rishavraj):

so \[y = \frac{ 12 }{ 13 }x - \frac{ 10 }{ 13 }\]

OpenStudy (anonymous):

Oh I understand now @rishavraj

rishavraj (rishavraj):

ok coolll

rishavraj (rishavraj):

@jadareese whts ur answer??

OpenStudy (anonymous):

C @rishavraj

rishavraj (rishavraj):

lol why???

rishavraj (rishavraj):

\[c = -\frac{ 10 }{ 13 }\]